OpenStudy (anonymous):

radical form of t ^-2/7

OpenStudy (anonymous):

the 2 in the numerator if the exponent means squared (power of two) the 7 in the denominator means the seventh root the minus sign means flip it (take the reciprocal) so \[\huge t^{-\frac{2}{7}}=\frac{1}{\sqrt[7]{t^2}}\]

OpenStudy (anonymous):

Thank you so much, @satellite73 !

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

gotta help anyone with a fender

OpenStudy (anonymous):

Haha, YES! @satellite73